OpenStudy (anonymous):

27. What is the equation, in slope- intercept form of the line that has a slope of 3/4 and passes through the point ( 4, -5)? A. y = 3/4x - 17 B. y = 3/4x - 2 C. y = 3/4x + 2 D. y = 3/4x - 8

OpenStudy (cj49):

y-y1=m(x-x1)

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

ty :) <3

OpenStudy (cj49):

sorry its D

OpenStudy (cj49):

y+5=3/4x-3 y=3/4x-8

OpenStudy (anonymous):

tyyy

OpenStudy (cj49):

np..c: